# ELMM: Efficient Lightweight Multimodal Large Language Models for Multimodal Knowledge Graph Completion

### v1

@@ Abstract
Multimodal Knowledge Graphs (MKGs) extend traditional knowledge graphs by incorporating visual and textual modalities, enabling richer and more expressive entity representations. However, existing MKGs often suffer from incompleteness, which hinder their effectiveness in downstream tasks. Therefore, multimodal knowledge graph completion (MKGC) task is receiving increasing attention. While large language models (LLMs) have shown promise for knowledge graph completion (KGC), their application to the multimodal setting remains underexplored. Moreover, applying Multimodal Large Language Models (MLLMs) to the task of MKGC introduces significant challenges: (1) the large number of image tokens per entity leads to semantic noise and modality conflicts, and (2) the high computational cost of processing large token inputs. To address these issues, we propose Efficient Lightweight Multimodal Large Language Models (ELMM) for MKGC. ELMM proposes a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism, which adaptively compresses image tokens from both textual and visual views, thereby effectively reducing redundancy while retaining necessary information and avoiding modality conflicts. Additionally, we design an attention pruning strategy to remove redundant attention layers from MLLMs, thereby significantly reducing the inference cost. We further introduce a linear projection to compensate for the performance degradation caused by pruning. Extensive experiments on benchmark FB15k-237-IMG and WN18-IMG demonstrate that ELMM achieves state-of-the-art performance while substantially improving computational efficiency, establishing a new paradigm for multimodal knowledge graph completion.
@@ 1 Introduction
Knowledge Graphs (KGs) represent real-world knowledge in a structured form using factual triples (head, relation, tail). In recent years, they have attracted considerable attention from both academia and industry, and have been widely applied to various downstream tasks such as information retrieval Yang (2020); Wang et al. (2024), event forecasting Huang et al. (2024); Gu and Krenn (2025), and recommendation systems Guo et al. (2020); Jiang et al. (2024). Multimodal Knowledge Graphs (MKGs) enhance traditional KGs by integrating information from multiple modalities, such as text and images, thereby enabling richer and more accurate knowledge representations. This multimodal fusion has shown to improve the performance of a wide range of intelligent systems Guo et al. (2024b). However, existing MKGs often suffer from incomplete information, which limits their effectiveness in practical applications. To address this challenge, the task of Multimodal Knowledge Graph Completion (MKGC) has emerged and has seen rapid development Gao et al. (2025); Zhang et al. (2024a); Chen et al. (2022); Shang et al. (2024). The core objective of MKGC is to leverage multimodal data to enhance the expressiveness of entity representations and uncover latent knowledge, thereby completing missing information in MKGs.
Large models, through extensive pre- and post-training on massive datasets, exhibit emergent capabilities and encode rich real-world knowledge Achiam et al. (2023); Guo et al. (2025a); Xiong et al. (2025), making them promising for knowledge graph completion (KGC). Recent studies have applied large language models (LLMs) to text-based KGC task. For example, KICGPT Wei et al. (2024) addresses the long-tail issue by re-ranking candidate triples from traditional methods, while MKGL Guo et al. (2024a) investigates whether LLMs can comprehend the structured representation of knowledge in the form of triples (head, relation, tail). However, current research largely overlooks multimodal inputs, leaving the potential of multimodal large language models(MLLMs) for MKGC underexplored.
In multimodal knowledge graphs (MKGs), entities are typically associated with various modalities of information, such as images and textual descriptions. For example, in datasets like FB15k-237-IMG Bollacker et al. (2008) and WN18-IMG Bordes et al. (2013), each entity is linked to 10 images along with corresponding textual descriptions, providing rich multimodal contextual information to complement the structured triples. State-of-the-art multimodal large language models (e.g., LLaVA-1.5Liu et al. (2024), Qwen2.5-VL Bai et al. (2025b), and MobileVLM Chu et al. (2023)) typically employ visual encoders, such as ViT Dosovitskiy et al. (2020), to encode different regions of an image into a sequence of image tokens. These image tokens are then concatenated with text tokens and jointly processed by MLLMs in a unified framework.
@@
However, directly applying the aforementioned multimodal fusion paradigms to the task of multimodal knowledge graph completion presents several critical challenges. First, the number of image tokens per entity is large and often redundant. Standard visual encoders generate hundreds of tokens per image; with 10 images per entity, this results in thousands of tokens. Moreover, entities may exhibit different semantics under varying relational contexts, making the relevance of individual image tokens highly context-dependent. As shown in Figure 1, visual regions of the entity Taylor Swift vary in importance across different relations. Naively concatenating all image and text tokens introduces noise and modality conflicts, hindering semantic alignment. Second, the computational cost of processing such large token inputs is prohibitive. In multimodal large language models, the computational complexity of attention mechanisms typically grows quadratically with the number of tokens, while memory consumption increases at least linearly. This limits the model’s applicability in real-time and large-scale MKGC scenarios.
@@
To address the aforementioned challenges, we propose ELMM, the first novel method to leverage Efficient Lightweight Multimodal Large Language Models (ELMM) for MKGC. First, we propose a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism to identify the most representative image tokens from both textual and visual views. This compressor adaptively compresses image tokens based on their semantic relevance to the textual description(containing entities and relations) and their intrinsic visual importance, compressing all image tokens from multiple input images into a fixed number of high-information tokens. This strategy significantly reduces visual noise and avoids modality conflicts, thereby enhancing both the inference efficiency and the representational capacity of the model. Second, through empirical observation, we identify a structural redundancy in the attention layers of MLLMs when applied to MKGC task—specifically, certain layers exhibit highly similar input and output representations, indicating limited contribution to overall performance. To mitigate this, we propose an attention pruning strategy that removes these redundant layers, effectively reducing computational overhead during inference. Additionally, we observe that the error introduced by the pruning strategy can be compensated by a linear projection, and in turn design a pipeline to initialize the linear projection.
Our contribution can be summarised as follows:
@@
  • We propose Efficient Lightweight Multimodal Large Language Models (ELMM) for multimodal knowledge graph completion, a novel MKGC method that designs a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism to adaptively compress image tokens from both textual and visual views. This mechanism enables effective compression of multimodal inputs while mitigating modality conflicts and reducing noise. To the best of our knowledge, this is the first work to address the MKGC task using Multimodal Large Language Models.
@@
  • We propose an attention pruning strategy that identifies and removes redundant attention layers in multimodal large language models (MLLMs) for the MKGC task, thereby reducing computational overhead. To mitigate pruning-induced errors, we apply linear projection as compensation.
@@
  • Extensive experiments on the FB15k-237-IMG and WN18-IMG datasets demonstrate the superiority of ELMM in the MKGC task.
@@
2 Related Work
2.1 Multimodal Knowledge Graph Completion
Multimodal knowledge graph completion (MKGC) seeks to enhance the reasoning ability of knowledge graphs by jointly leveraging structural triples and multimodal entity information such as images and text. Early approaches Wang et al. (2019, 2021); Xie et al. (2016) typically used simple feature concatenation, but recent research has introduced more advanced techniques for fine-grained and adaptive multimodal fusion. Representative models include MKGformer Chen et al. (2022), which employs hybrid transformers with multi-level fusion; SGMPT Liang et al. (2024) introduces structural information based on MKGformer; LAFA Shang et al. (2024), which adopts link-aware fusion and neighbor aggregation; MyGO Zhang et al. (2024a), which tokenizes multimodal entity information for fine-grained representation; and NativE Zhang et al. (2024b), which addresses modality imbalance in real-world scenarios.
With the rapid advancement of multimodal large language models (MLLMs), these models have demonstrated substantial potential in task related to Multimodal Knowledge Graphs (MKGs). Although existing approaches have achieved notable progress in this area, most studies have yet to fully explore and leverage the capabilities of MLLMs. To address this gap, we propose ELMM, a novel method designed to harness the strengths of MLLMs.
2.2 Multimodal Large Language Models
With the rapid development of Multimodal Large Language Models(MLLMs), such as GPT-4 Achiam et al. (2023), LLaVA-1.5 Liu et al. (2024), Qwen2.5-VL Bai et al. (2025b), and Emu2 Sun et al. (2024), vision and language understanding have been deeply integrated within a unified semantic space. These models align visual and textual inputs through large-scale pretraining, enabling strong performance in visual question answering, captioning, and knowledge reasoning. Recent advances also focus on knowledge-enhanced pretraining Ma et al. (2024), multi-modal adapters Duan et al. (2024); Chen et al. (2025); Seputis et al. (2024), and efficient token selection to further improve performance in knowledge-intensive scenarios.
Despite the remarkable progress made by MLLMs across various multimodal tasks, they have yet to demonstrate significant breakthroughs in the task of MKGC. To address this limitation, we propose ELMM that incorporates several innovative mechanisms, including a Multi-view Visual Token Compressor (MVTC), an attention pruning strategy, and a multimodal knowledge reasoning completion layer. These components collectively enhance the capability of MLLMs in the MKGC task, leading to state-of-the-art performance in this domain for the first time.
@@ 3.1.1 Problem Formulation
A multimodal knowledge graph (MKG) is defined as a directed graph $MKG=(\mathcal{E},\mathcal{R},\mathcal{G},\mathcal{A}_{\mathcal{M}})$, where $\mathcal{E}$ and $\mathcal{R}$ denote the sets of entities and relations, respectively. The factual triples are represented as $\mathcal{G}=\{(e_{h},r_{h,t},e_{t})\mid e_{h},e_{t}\in\mathcal{E},r_{h,t}\in\mathcal{R}\}$. Each entity is associated with a set of multimodal attributes $\mathcal{A}_{\mathcal{M}}$, which includes textual $\mathcal{M}_{t}$ and visual $\mathcal{M}_{v}$ modalities. The task of multimodal knowledge graph completion (MKGC) aims to infer missing entities in $\mathcal{G}$. Given an incomplete triple $(h,r,?)$ or $(?,r,t)$, the goal is to predict the missing tail entity $t$ or $h$.
@@ 3.4 Attention Pruning Strategy
Due to the high computational cost of attention mechanisms, they often introduce significant inference latency. For instance, with an input sequence length of 512 tokens, the attention module alone can introduce up to 0.3437 seconds of latency per inference, accounting for approximately 77.9% of the total inference time (0.4410 seconds). To mitigate this overhead, we perform a redundancy analysis of the attention modules in MLLMs, such as LLaVA-1.5 Liu et al. (2024), during MKGC task. Specifically, we measure the cosine similarity between the input and output of each attention module to assess its redundancy. As illustrated in Figure 3, the upper layers of the model consistently exhibit high input-output similarity across various experimental settings. This observation suggests a significant degree of computational redundancy in certain attention module of MLLMs during MKGC task.
@@ 4 Experiments
ELMM(Ours) 112 34.1 47.3 60.7 11 95.7 97.5 98.7
{*} indicates that the result comes from SGMPT Liang et al. (2024).
@@ 4.1 Datasets
We evaluate the performance of ELMM on two publicly available multimodal knowledge graph datasets: FB15k-237-IMG Bollacker et al. (2008) and WN18-IMG Bordes et al. (2013). Each dataset comprises three modalities: (1) structured knowledge graph triples, (2) textual descriptions of entities, and (3) a set of 10 associated images per entity. Detailed statistics for both datasets are presented in Table 1.
@@ 4.2 Setting
We adopt LLaVA-1.5 7B Liu et al. (2024) as the base multimodal large language model (MLLM), incorporating Low-Rank Adaptation (LoRA) Hu et al. (2022) techniques into both the query and value layers to enable parameter-efficient fine-tuning. All experiments are conducted using eight NVIDIA A100 GPUs, and both training and inference are implemented with the PyTorch framework. The parameter $K$ in the attention pruning strategy is always set to 16, indicating that 16 attention modules in the MLLM are pruned and replaced with linear projections.
@@
For a comprehensive performance comparison, we benchmark our model against two representative categories of methods: (i) multimodal knowledge graph completion methods, and (ii) uni-modal knowledge graph completion methods based on large language models (LLMs). The former category includes IKRL Xie et al. (2016), TransAE Wang et al. (2019), RSME Wang et al. (2021), KG-BERT Yao et al. (2019), VisualBERT Li et al. (2019), ViLBERT Su et al. (2019), MKGformer Chen et al. (2022), LAFA Shang et al. (2024), NativE Zhang et al. (2024b), SGMPT Liang et al. (2024), MyGO Zhang et al. (2024a), MPIKGC Xu et al. (2024), and AdaMF-MAT Zhang et al. (2024c). The LLM-based uni-modal knowledge graph completion baselines include KICGPT Wei et al. (2024), K-ON Guo et al. (2025b) and MKGL Guo et al. (2024a). the traditional most powerful multimodal methods
@@ 4.3 Comparison with State-of-the-art
We conduct a comprehensive comparison between the proposed ELMM method and sixteen state-of-the-art models across two benchmark datasets. As shown in Table 2, ELMM consistently outperforms all competing methods across all evaluation metrics. Notably, on the FB15k-237-IMG dataset, ELMM demonstrates substantial improvements over most powerful multimodal approaches, achieving relative gains of 26.8%, 18.8%, and 10.2% in Hits@1, Hits@3, and Hits@10, respectively. These results strongly validate the powerful potential of multimodal large language models (MLLMs) in the task of multimodal knowledge graph completion. Furthermore, compared to uni-modal methods based on large language models, ELMM exhibits clear advantages on both datasets, underscoring its effectiveness and robustness in extracting and integrating multimodal information.
@@ 4.4 Ablation Study
Notably, ELMM achieves over 30% faster inference without significant performance degradation compared to the non-pruned model, indicating redundancy in attention modules for the MKGC task. Moreover, removing the linear compensation or initializing its projection weights $W_{c}$ to zero both lead to comparable performance drops, suggesting that the effectiveness of linear compensation depends critically on appropriate initialization.
@@
ELMM 112 34.1 47.3 60.7
- w/o Image 122 33.7 47.0 60.2
- w/o Text 156 30.8 43.5 57.8
- w/o MVTC 193 27.9 39.8 54.3
- w/o Pruning 106 33.8 47.2 61.0
- w/o Linear 159 33.2 44.9 58.2
- w/o Init 147 33.5 44.8 58.9
- Head Layer 121 33.9 46.9 60.6

### v2

@@ Abstract
Multimodal Knowledge Graphs (MKGs) extend traditional knowledge graphs by incorporating visual and textual modalities, enabling richer and more expressive entity representations. However, existing MKGs often suffer from incompleteness, which hinders their effectiveness in downstream tasks. Therefore, multimodal knowledge graph completion (MKGC) task is receiving increasing attention. While large language models (LLMs) have shown promise for knowledge graph completion (KGC), their application to the multimodal setting remains underexplored. Moreover, applying Multimodal Large Language Models (MLLMs) to the task of MKGC introduces significant challenges: (1) the large number of image tokens per entity leads to semantic noise and modality conflicts, and (2) the high computational cost of processing large token inputs. To address these issues, we propose Efficient Lightweight Multimodal Large Language Models (ELMM) for MKGC. ELMM proposes a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism, which adaptively compresses image tokens from both textual and visual views, thereby effectively reducing redundancy while retaining necessary information and avoiding modality conflicts. Additionally, we propose an attention pruning strategy to remove redundant attention layers in MLLMs, substantially reducing inference cost. To mitigate the performance degradation introduced by pruning, we further incorporate a linear projection for error compensation. Extensive experiments on four benchmark datasets demonstrate that ELMM achieves state-of-the-art performance.
@@ 1 Introduction
Knowledge Graphs (KGs) represent real-world knowledge in a structured form using factual triples (head, relation, tail). In recent years, they have attracted considerable attention from both academia and industry, and have been widely applied to various downstream tasks such as information retrieval 2020Biomedical; wang2024knowledge, and recommendation systems guo2020survey; jiang2024diffkg. Multimodal Knowledge Graphs (MKGs) enhance traditional KGs by integrating information from multiple modalities, such as text and images, thereby enabling richer and more accurate knowledge representations. This multimodal fusion has shown to improve the performance of a wide range of intelligent systems guo2024lgmrec. However, existing MKGs often suffer from incomplete information, which limits their effectiveness in practical applications. To address this challenge, the task of Multimodal Knowledge Graph Completion (MKGC) has emerged and has seen rapid development gao2025mixed; MyGO; MKGformer; LAFA. The core objective of MKGC is to leverage multimodal data to enhance the expressiveness of entity representations and uncover latent knowledge, thereby completing missing information in MKGs.
Large models, through extensive pre- and post-training on massive datasets, exhibit emergent capabilities and encode rich real-world knowledge guo2025deepseek; xiong2025uniattn, making them promising for KGC. For example, KICGPT Kicgpt addresses the long-tail issue by re-ranking candidate triples from traditional methods, while MKGL mkgl investigates whether LLMs can comprehend the structured representation of knowledge in the form of triples (head, relation, tail). However, current research largely overlooks multimodal inputs, leaving the potential of multimodal large language models (MLLMs) for MKGC underexplored.
In MKGs, entities are typically associated with various modalities of information, such as images and textual descriptions. For example, in datasets like FB15k-237-IMG FB15K-237 and WN18-IMG WN18, each entity is linked to 10 images along with corresponding textual descriptions, providing rich multimodal contextual information to complement the structured triples. State-of-the-art MLLMs (e.g., LLaVA-1.5llava1.5, Qwen2.5-VL qwen2.5vl, and MobileVLM chu2023mobilevlm) typically employ visual encoders, such as ViT vit, to encode different regions of an image into a sequence of image tokens. These image tokens are then concatenated with text tokens and jointly processed by MLLMs in a unified framework.
@@
However, directly applying existing multimodal fusion paradigms to MKGC poses several challenges. First, the number of image tokens per entity is large and often redundant. Standard visual encoders generate hundreds of tokens per image; with 10 images per entity, this results in thousands of tokens. Moreover, the semantic relevance of visual regions is relation-dependent: as illustrated in Figure 1, different relations emphasize different regions of Taylor Swift. Naively concatenating all visual and textual tokens therefore introduces noise and cross-modal interference, impeding effective semantic alignment. Second, the computational cost of processing such large token inputs is prohibitive. In MLLMs, the computational complexity of attention mechanisms typically grows quadratically with the number of tokens, while memory consumption increases at least linearly. This limits the model’s applicability in large-scale MKGC scenarios.
@@
To address the above challenges, we propose ELMM, the first novel method to leverage Efficient Lightweight Multimodal Large Language Models (ELMM) for MKGC. First, we propose a Multi-view Visual Token Compressor (MVTC) based on multi-head attention to select the most informative visual tokens from both textual and visual views. MVTC adaptively compresses visual tokens according to their semantic relevance to the textual description (entities and relations) and their intrinsic visual salience, aggregating tokens from multiple images into a fixed set of high-information representations. This design effectively reduces visual noise and modality interference, improving both inference efficiency and representational capacity. Second, we empirically identify structural redundancy in MLLM attention layers for MKGC, where certain layers produce highly similar input–output representations and contribute marginally to performance. To address this, we propose an attention pruning strategy that removes such redundant layers, significantly reducing inference overhead. We further observe that the pruning-induced error can be effectively compensated by a linear projection, and accordingly design a principled initialization pipeline for this projection.
Our contributions can be summarised as follows:
@@
  • We propose Efficient Lightweight Multimodal Large Language Models (ELMM) for multimodal knowledge graph completion, a novel MKGC method that designs a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism to adaptively compress image tokens from both textual and visual views. This mechanism enables effective compression of multimodal inputs while mitigating modality conflicts and reducing noise. To the best of our knowledge, this is the first work to address the MKGC task using MLLMs.
@@
  • We propose an attention pruning strategy that identifies and removes redundant attention layers in MLLMs for the MKGC task, thereby reducing computational overhead. To mitigate pruning-induced errors, we apply linear projection as compensation.
@@
  • Extensive experiments on four benchmark datasets demonstrate the superiority of ELMM in the MKGC task.
@@ 2.1.1 Problem Formulation
A multimodal knowledge graph (MKG) is defined as a directed graph $MKG=(\mathcal{E},\mathcal{R},\mathcal{G},\mathcal{A}_{\mathcal{M}})$, where $\mathcal{E}$ and $\mathcal{R}$ denote the sets of entities and relations, respectively. The factual triples are represented as $\mathcal{G}=\{(e_{h},r_{h,t},e_{t})\mid e_{h},e_{t}\in\mathcal{E},r_{h,t}\in\mathcal{R}\}$. Each entity is associated with a set of multimodal attributes $\mathcal{A}_{\mathcal{M}}$, which includes textual $\mathcal{M}_{t}$ and visual $\mathcal{M}_{v}$ modalities. The task of multimodal knowledge graph completion (MKGC) aims to infer missing entities in $\mathcal{G}$. Given an incomplete triple $(h,r,?)$ or $(?,r,t)$, the goal is to predict the missing entity $t$ or $h$.
@@ 2.4 Attention Pruning Strategy
Due to the high computational cost of attention mechanisms, they often introduce significant inference latency. For instance, with an input sequence length of 512 tokens, the attention module alone can introduce up to 0.3437 seconds of latency per inference, accounting for approximately 77.9% of the total inference time (0.4410 seconds). To mitigate this overhead, we perform a redundancy analysis of the attention modules in MLLMs, such as LLaVA-1.5 llava1.5, during MKGC task. Specifically, we measure the cosine similarity between the input and output of each attention module to assess its redundancy. As illustrated in Figure 3, the upper layers of the model consistently exhibit high input-output similarity across various experimental settings. This observation suggests a significant degree of computational redundancy in certain attention modules of MLLMs during MKGC task.
@@ 3 Experiments
GLTW (ACL 2025) - 35.1 48.1 61.4 17 95.2 96.3 97.9
PEKGC (EMNLP 2025) - 33.6 46.6 54.4 30 94.8 96.0 96.9
ELMM(Ours) 105 37.4 50.2 63.7 11 96.1 97.8 98.9
{*} indicates that the result comes from SGMPT SGMPT.
@@ 3.1 Datasets
We evaluate the performance of ELMM on four publicly available multimodal knowledge graph datasets: FB15k-237-IMG FB15K-237, WN18-IMG WN18, DB15KDB15K and MKG-WMMRNS. Each dataset comprises three modalities: (1) structured knowledge graph triples, (2) textual descriptions of entities, and (3) a set of associated images per entity. Detailed statistics for both datasets are presented in Table 1. Due to space constraints, we report evaluation results primarily on the FB15k-237-IMG and WN18-IMG datasets in the main text; results on DB15K and MKG-W are provided in Appendix D.
@@ 3.2 Setting
We adopt LLaVA-1.5 7B llava1.5 as the base multimodal large language model (MLLM), incorporating Low-Rank Adaptation (LoRA) hu2022lora techniques into both the query and value layers to enable parameter-efficient fine-tuning. The parameter $K$ in the attention pruning strategy is always set to 16, indicating that 16 attention modules in the MLLM are pruned and replaced with linear projections. Full hyperparameter configurations are provided in the Appendix C.
@@
For a comprehensive performance comparison on FB15k-237-IMG and WN18-IMG datasets, we benchmark our model against two representative categories of methods: (i) multimodal knowledge graph completion methods, and (ii) uni-modal knowledge graph completion methods. The former category includes TransAE TransAE, RSME RSME, KG-BERT kg-bert, VisualBERT visualbert, ViLBERT Vilbert, MKGformer MKGformer, LAFA LAFA, NativE zhang2024native, SGMPT SGMPT, MyGO MyGO, MPIKGC MPIKGC, and AdaMF-MAT AdaMF-MAT. The LLM-based uni-modal knowledge graph completion baselines include KICGPT Kicgpt, K-ON kno, GLTW GLTW, PEKGC SAT and MKGL mkgl.
@@ 3.3 Comparison with State-of-the-art
We conduct a comprehensive comparison between the proposed ELMM method and 18 state-of-the-art models across FB15k-237-IMG and WN18-IMG datasets. As shown in Table 2, ELMM consistently outperforms all competing methods across all evaluation metrics. Notably, on the FB15k-237-IMG dataset, ELMM demonstrates substantial improvements over most powerful multimodal approaches, achieving relative gains of 39.0%, 26.1%, and 15.6% in Hits@1, Hits@3, and Hits@10, respectively. These improvements primarily stem from ELMM’s ability to directly leverage the expressive reasoning capabilities of MLLMs to jointly model multimodal information within a unified semantic space. As a result, ELMM more effectively resolves ambiguous cases and entities that require fine-grained visual reasoning. These results strongly demonstrate the potential of MLLMs for MKGC. Furthermore, compared to uni-modal methods based on large language models, ELMM exhibits clear advantages on both datasets, underscoring its effectiveness and robustness in extracting and integrating multimodal information.
@@ 3.4 Ablation Study
Notably, ELMM achieves over 30% faster inference without significant performance degradation compared to the non-pruned model, indicating redundancy in attention modules for the MKGC task. Moreover, removing the linear compensation or initializing its projection weights $W_{c}$ to zero both lead to comparable performance drops, suggesting that the effectiveness of linear compensation depends critically on appropriate initialization. Also, removing the multimodal knowledge reasoning completion layer similarly hurts overall performance.
@@
ELMM 105 37.4 50.2 63.7
- w/o Image 115 37.6 49.8 63.1
- w/o Text 146 33.5 46.3 60.7
- w/o MVTC 187 28.5 41.5 56.2
- w/o Pruning 102 37.1 50.0 64.2
- w/o Linear 146 35.9 47.7 60.6
- w/o Init 140 36.8 47.5 61.5
- Head Layer 119 37.2 49.2 63.0